OpenStudy (anonymous):

Choose the Standard Form of the equation y = − 1/3x - 4. A) 1/3x + y = −4 B) 3x + y = −4 C) x + 3y = −12 D) −3x + y = −12

OpenStudy (anonymous):

The standard form is ax+by=c, where 'a' must be a positive whole number. So you first set the equation in the standard form leaving: 1/3x+y=-4. Then, you multiply the equation by the denominator to make 'a' be a a whole number. 3(1/3x+y=-4). Finally leaving: x+3y=-12 -or- C.